A capacitor has parallel plates that have an area of 1.09 cm2 and are 1.08 mm apart. There is a vacuum between the plates, and the electric field magnitude is 1.10 106 N/C. (a) Find the potential across the plates. (b) Find the charge.
Answer #1

a] Potential = Ed = 1.1e6 V/m * 1.08e-3 m = 1188 V answer

b] Charge = Area*sigma = area*E*e0 = 1.09e-4*1.1e6*8.85e-12

= 1.06*10^-9 C answer